OpenStudy (anonymous):

10.02] Which ratio forms a proportion with 6 over 14.? 12 over 7. 12 over 28. 7 over 3. 4 over 12.

OpenStudy (anonymous):

6/14=3/7=12/28

OpenStudy (anonymous):

rylee

OpenStudy (anonymous):

i have six more

OpenStudy (anonymous):

okay but how did u get that????